OpenStudy (anonymous):

*Advance Algebra With Financial Applications* Kirby made contributions to a Roth IRA over the course of 35 working years. His contributions averaged $2,000 annually. Kirby was in the 26% tax bracket during his working years. The average annual rate of return on the account was 5%. Upon retirement, Kirby stopped working and making Roth IRA contributions. Instead, he started living on withdrawals from the retirement account. At this point, Kirby dropped into the 15% tax bracket. Factoring in taxes, what is the effective value of Kirby’s Roth IRA at retirement? Assume annual compounding.

OpenStudy (anonymous):

$170,140.61 $172,940.61 $162,440.61 $133,674.05

OpenStudy (amistre64):

this one is tricky, or at least the one that was like it

OpenStudy (anonymous):

I was think of just 2000*26%*5%*15% = ?

OpenStudy (amistre64):

A roth IRA is such that the interest earned is not taxed; and the contributions to it are taxed prior to making the contribution.... but then i might have a different view of a roth ira then your material

OpenStudy (amistre64):

as such, my concept would amount to a 2000 annuity for 35 years

OpenStudy (amistre64):

also, had an issue with the term "effective value"

OpenStudy (anonymous):

FV = PV(1+i/n)^(nt)

OpenStudy (anonymous):

FVOA=2000*(((1+0.05))^1.35-1)/(0.05))

OpenStudy (amistre64):

FVOA-D is more like it .... contributes 2000, before years up ... then applied

OpenStudy (anonymous):

Is my formula right ?

OpenStudy (amistre64):

theres a typo, but the structure seems valid using an ordinary annuity setup

OpenStudy (amistre64):

^35 ... not ^1.35

OpenStudy (anonymous):

Tax=26 % of 2000 * 35

OpenStudy (amistre64):

the contributions are already taxed .... they dont get taxed again

OpenStudy (amistre64):

when i get a paycheck, i dont have to pay taxes on the after tax amount that I get to put into my pocket

OpenStudy (anonymous):

0.26 * 2000 * 35

OpenStudy (amistre64):

you traditionally get taxed for income that you earn .... not income that you have already earned and already been taxed for

OpenStudy (anonymous):

520 * 35

OpenStudy (anonymous):

18,200

OpenStudy (anonymous):

he would pay that in income

OpenStudy (anonymous):

now the final value we find

OpenStudy (anonymous):

180640.6147 - 18,200 = ?

OpenStudy (anonymous):

it gave me 162,440.61<-------- :D OMG !! OMG !! I DID IT LOL OMG ! I GOT THE ANSWER ! @amistre64

OpenStudy (amistre64):

good job then :) still makes no sense from my end tho